OpenStudy (anonymous):

can you help proof this equation??? 5u-u=u+15

OpenStudy (anonymous):

u must equal 5,using basic algebraic manipulation

OpenStudy (anonymous):

subtract 5u-u to get 4u, then you subtract u from u+15 to get 3u=15. divide by 3

OpenStudy (anonymous):

5u-u=u+15 4u=u+15 4u-u=15 3u=15 u=15/3 u=5 now put this value into eq:

OpenStudy (anonymous):

eeyup